# Counting Distinct Elements in the Turnstile Model with Differential Privacy under Continual Observation

### v1

@@ 4.2 From the reduction to the accuracy lower bound
We use the following lower bound on inner product queries. Like a similar lower bound of Mir et al. [49], it uses the reconstruction attacks of [21, 24], together with the argument of De [18] that rules out reconstruction from the outputs of $(\varepsilon,\delta)$-differentially private algorithms.
Theorem 4.5 (Inner product queries lower bound (based on [21, 24, 49, 18]).
There are constants $c_{1},c_{2}>0$ such that, for sufficiently large $n>0$: if $\mathcal{A}$ is $\alpha$-accurate for $\mathsf{InnerProducts}_{k,n}$ (Definition 4.1) with $k=c_{1}n$ and $\alpha=c_{2}\sqrt{n}$, then $\mathcal{A}$ is not $(1,\frac{1}{3})$-differentially private.
We first prove Theorem 1.5 for $\varepsilon=1$ and then boost it to arbitrary $\varepsilon<1$ using the reduction in Theorem B.1.
@@
Let $\delta=o\left(\frac{1}{T}\right)$ and sufficiently large $w,T\in\mathbb{N}$ such that $w\leq T$. If there exists a $(1,\delta)$-event-level-DP mechanism that is $\alpha$-accurate for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$, then $\alpha=\Omega(\mathsf{min}(\sqrt{w},T^{1/4}))$.
@@
Fix sufficiently large $w$ such that $w\leq\sqrt{T}$. Let $c_{1},c_{2}>0$ be the constants from Theorem 4.5. Assume that $\mathcal{M}$ is a $\left(1,o(\frac{1}{T})\right)$-event-level-DP, $(\frac{c_{2}}{2}\sqrt{w})$accurate mechanism for $\mathsf{CountDistinct}$ for turnstile streams of length $T$ with maximum flippancy at most $w$. Then, set $k=\frac{w}{2}$ and $n=\frac{k}{c_{1}}=\frac{w}{2c_{1}}$. This choice of $k$ and $n$ satisfies the conditions of Lemma 4.2 since the flippancy of the stream is at most $w=2k$ and for $w\leq\sqrt{T}$ we have that $(2k+1)n=(w+1)\frac{w}{2c_{1}}\leq w^{2}\leq T$. Therefore, $\mathcal{A}$ (Algorithm 4) is $\left(1,o(\frac{1}{T})\right)$-DP and $(c_{2}\sqrt{w})$-accurate for $\mathsf{InnerProducts}_{k,n}$. Since $\frac{1}{T}\leq\frac{1}{n}$ and $w=O(n)$, we get that $\mathcal{A}$ is $(1,o(\frac{1}{n}))$-DP and $c_{2}\sqrt{n}$-accurate for $\mathsf{InnerProducts}_{k,n}$, where $k=c_{1}n$.
@@
Finally, we invoke the reduction in Theorem B.1 to improve the dependence on $\varepsilon$ and complete the proof of Theorem 1.5.
@@
First, for $\varepsilon<\frac{2}{T}$ we obtain an error lower bound of $\Omega(T)$, via a group privacy argument that is exactly the same as in the item-level lower bound (we direct the reader to the proof of Theorem 1.6 for more details). Next, consider the case where $\varepsilon\geq\frac{2}{T}$. Lemma 4.6 provides a lower bound of $\alpha^{\prime}=\Omega\Big{(}\mathsf{min}\left(\sqrt{w},T^{1/4}\right)\Big{)}$ for ($\varepsilon^{\prime}=1$,$\delta^{\prime}=o(1/T)$)-event-level-DP, $\alpha^{\prime}$-accurate mechanisms for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$. By invoking Theorem B.1, we obtain the following lower bound on accuracy for $(\varepsilon,\delta)$-DP mechanisms where $\delta=\frac{\delta^{\prime}\varepsilon}{10}=o(\frac{\varepsilon}{T})$:
@@ 5 Item-level privacy lower bound
In this section, we prove Theorem 1.6 that provides strong lower bounds on the accuracy parameter $\alpha$ of any item-level differentially private mechanism for $\mathsf{CountDistinct}$ in the continual release model for turnstile streams. This lower bound is parameterized by $w$, the maximum flippancy of the input stream.
@@ 5.1 Reduction from $\mathrm{\sf Marginals}$
To prove our lower bounds for $\mathsf{CountDistinct}$, we reduce from the problem of approximating 1-way marginals in the batch model. The reduction is presented in Algorithm 5. The privacy and accuracy guarantees of our reduction are stated in Lemma 5.2. In Section 5.2, we use Lemma 5.2 to complete the proof of Theorem 1.6.
We first provide a high-level overview of our reduction. The function $\mathrm{\sf Marginals}_{n,d}:\{0,1\}^{n\times d}\to[0,1]^{d}$ maps a dataset $y$ of $n$ records and $d$ attributes to a vector $(q_{1}(y),\dots,q_{d}(y)),$ where $q_{j}$, called the $j^{th}$ marginal, is defined as $q_{j}(y)=\frac{1}{n}\sum_{i=1}^{n}y[i][j].$ Let $\mathcal{M}$ be an $(\varepsilon,\delta)$-DP and $\alpha$-accurate mechanism for $\mathsf{CountDistinct}$ in the continual release model. We use $\mathcal{M}$ to construct an $(\varepsilon,\delta)$-DP batch algorithm $\mathcal{A}$ that is $(\frac{\alpha}{n})$-accurate for $\mathrm{\sf Marginals}_{n,d}$. Consider a universe $\mathcal{U}=[n]\cup\{\perp\}$ for $\mathsf{CountDistinct}:\mathcal{U}_{\pm}^{T}\to\mathbb{N}$. The main idea in the construction (presented in Algorithm 5) is to force $\mathcal{M}$ to output an estimate of the marginals, one attribute at a time. Given a dataset $y\in\{0,1\}^{n\times d}$, to get an estimate of the first marginal, $\mathcal{A}$ sends element $i$ to $\mathcal{M}$ for each record $y[i]$ with a $1$ in the first attribute. We call this phase one of estimating the first marginal. The answer produced by $\mathcal{M}$ at this point is an estimate of the sum of the first attribute of all records $y[1],\dots,y[n]$. This can be divided by $n$ to estimate the first marginal. Then, to ‘clear the slate’, $\mathcal{A}$ sends $-i$ to $\mathcal{M}$ for each $y[i]$ with a $1$ in the first attribute. We call this phase two of estimating the first marginal. It repeats this for each attribute, collecting the answers from $\mathcal{M},$ and then outputs its estimates for the marginals. In actuality, in both phase one and phase two of estimating the $j^{\text{th}}$ marginal, $\mathcal{A}$ will input $\perp$ for each $y[i]$ that has a 00 in the $j^{\text{th}}$ attribute. This algorithm is $(\varepsilon,\delta)$-DP for $\mathrm{\sf Marginals}_{n,d}$ since changing one record $y[i]$ in the input to the algorithm $\mathcal{A}$ will only add or remove occurrences of a single element $i$ in the input to the mechanism $\mathcal{M}$.
@@
Let $\mathcal{A}$ be Algorithm 5. For all $\varepsilon>0,\delta\geq 0,\alpha\in\mathbb{R}^{+}$ and $d,n,w,T\in\mathbb{N}$, where $T\geq 2dn$ and $w\geq 2d$, if mechanism $\mathcal{M}$ is $(\varepsilon,\delta)$-item-level-DP and $\alpha$-accurate for $\mathsf{CountDistinct}$ for streams of length $T$ with maximum flippancy at most $w$ in the continual release model, then batch algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-DP and $\frac{\alpha}{n}$-accurate for $\mathrm{\sf Marginals}_{n,d}$.
@@
We start by reasoning about privacy. Fix neighboring datasets $y$ and $y^{\prime}$ that are inputs to batch algorithm $\mathcal{A}$ (Algorithm 5). (Datasets $y$ and $y^{\prime}$ differ in one row.) Let $x$ and $x^{\prime}$ be the streams constructed in Step 8 of $\mathcal{A}$ when it is run on $y$ and $y^{\prime}$, respectively. By construction, $x$ and $x^{\prime}$ are item-neighbors. Since $\mathcal{M}$ is $(\varepsilon,\delta)$-item-level-DP, and $\mathcal{A}$ only post-processes the outputs received from $\mathcal{M}$, closure under postprocessing (Lemma 2.1) implies that $\mathcal{A}$ is $(\varepsilon,\delta)$-DP.
@@
Notice that: (1) The coins of $\mathcal{A}$ are the same as the coins of $\mathcal{M}$ (since the transformation from $\mathcal{M}$ to $\mathcal{A}$ is deterministic). (2) The marginals computed in Step 10 of Algorithm 5 are computed using the relationship described by Equation (7). (3) The maximum flippancy of the stream constructed in Algorithm 5 is at most $2d$, since each item $i\in\mathcal{U}$ is added and removed at most once in each $z^{(j)}$ for $j\in[d]$. We obtain that $\mathcal{A}$ inherits its probability of success from $\mathcal{M}$:
@@ 5.2 From the reduction to the accuracy lower bound
For all $\varepsilon\in(0,1]$, $\delta\in[0,1]$, $\gamma\in(0,1)$, $d,n\in\mathbb{N}$, and algorithms that are $(\varepsilon,\delta)$-differentially private and $\gamma$-accurate for $\mathrm{\sf Marginals}_{n,d}$, the following statements hold.
@@
To prove Theorem 1.6, we proceed in two parts. First we show that the lower bound holds for $\varepsilon=1$. Then we use Theorem B.1 to extend it to all $\varepsilon<1$. Recall that the approximate-DP lower bound (on the error term $\alpha$) in Theorem 1.6 is the minimum of two terms. To prove this bound, we need to establish that, for every possible range of parameters, at least one term serves as a lower bound for $\alpha$.
@@
Let $\delta=o\left(\frac{1}{T}\right)$, and sufficiently large $w,T\in\mathbb{N}$ such that $w\leq T$. If there exists an $(1,\delta)$-item-level-DP mechanism that is $\alpha$-accurate for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$, then
$\alpha=\Omega\Big{(}\mathsf{min}\left(\frac{\sqrt{w}}{\log w},\frac{T^{1/3}}{\log T}\right)\Big{)}\text{ for approximate DP and }\alpha=\Omega\Big{(}\mathsf{min}\left(w,\sqrt{T}\right)\Big{)}\text{ when }\delta=0.$
@@
Let $\mathcal{A}$ be the algorithm for $\mathrm{\sf Marginals}_{n,d}$ with black-box access to an $\alpha$-accurate mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$, as defined in Algorithm 5. If $T\geq 2dn$ and $w\geq 2d$, then by Lemma 5.2, algorithm $\mathcal{A}$ is $(1,\delta)$-differentially private and $(\frac{\alpha}{n})$-accurate for $\mathrm{\sf Marginals}_{n,d}$. We can then use Lemma 5.3 to lower bound $\alpha$.
Approximate DP:
@@
Now suppose $w=T^{2/3}$. The above argument gives a lower bound of $\Omega\left(\frac{\sqrt{T^{2/3}}}{\log T^{2/3}}\right)$ on the accuracy of $\mathcal{M}$, and this lower bound applies to all $w>T^{2/3}$, since a mechanism that is $\alpha$-accurate for streams with maximum flippancy at most $w>w^{\prime}$ is also $\alpha$-accurate for streams with maximum flippancy at most $w^{\prime}$.
Pure DP:
The proof for when $\delta=0$ proceeds along the same lines, except that we consider the cases $w\leq\sqrt{T}$ and $w>\sqrt{T}$ and use Item 2 from Lemma 5.3 instead of Item 1: Suppose $w\leq\sqrt{T}$. Pick a dimension $d=w/2$, and number of entries $n=\frac{T}{w}$. If $\frac{\alpha}{n}<1$, then by Lemma 5.2 and Item 1 of Lemma 5.3, $n=\Omega\left(\frac{n\cdot d}{\alpha\cdot\varepsilon}\right)$ which means that $\alpha=\Omega\left(\frac{d}{\varepsilon}\right)=\Omega(w)$. Otherwise, if $\alpha\geq n$, then $\alpha\geq\frac{T}{w}\geq\sqrt{T}\geq w$.
@@
This concludes the proof of Lemma 5.4
Finally, we can extend the lower bounds for $\varepsilon=1$ from Lemma 5.4 to the general case of $\varepsilon<1$ using Theorem B.1.
@@
Suppose $\varepsilon<\frac{2}{T}$. For these values of $\varepsilon$, we prove an error lower bound of $\Omega(T)$, via a group privacy argument. Suppose for the sake of contradiction that $\alpha\leq T/4$. Consider universe $\mathcal{U}=[T]$. Let $x=\bot^{T}$ and $x^{\prime}$ be a stream of length $T$ such that $x[t]=t$ for all $t\in[T]$. These data streams differ on $T$ items. Let $r[T]$ and $r^{\prime}[T]$ be the final outputs of $\mathcal{M}$ on input streams $x$ and $x^{\prime}$, respectively. By the accuracy of $\mathcal{M}$, we have $\Pr[r[T]\leq T/4]\geq 0.99.$ Applying Lemma 2.3 on group privacy with $\varepsilon\leq{2}/{T}$ and group size $\ell=T$, we get $\Pr[r^{\prime}[T]>T/4]\leq e^{2}\cdot\Pr[r[T]>T/4]+\frac{2\delta}{\varepsilon}\leq e^{2}\cdot 0.01+o(\frac{1}{T})<0.99$ for sufficiently large $T$. But $\mathsf{CountDistinct}(x^{\prime})=T$, so $\mathcal{M}$ is not $T/4$-accurate for $x^{\prime}$, a contradiction. Hence, $\alpha=\Omega(T)$.
@@
Now suppose $\varepsilon\geq\frac{2}{T}$. For $\delta>0$, Lemma 5.4 provides a lower bound of $\alpha^{\prime}=\Omega\Big{(}\mathsf{min}\left(\frac{\sqrt{w}}{\log w},\frac{T^{1/3}}{\log T}\right)\Big{)}$ on accuracy for for ($\varepsilon^{\prime}=1$,$\delta^{\prime}=o(1/T)$)-item-level-DP, $\alpha^{\prime}$-accurate mechanisms for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$. By invoking Theorem B.1, we can extend this to the following lower bound for $(\varepsilon,\delta)$-DP mechanisms where $\delta=\frac{\delta^{\prime}\varepsilon}{10}=o(\frac{\varepsilon}{T})$:
@@
$\alpha=\frac{1}{\varepsilon}\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\log w},\frac{(\varepsilon T)^{1/3}}{\log\varepsilon T}\right)\right)=\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon\log w},\frac{T^{1/3}}{\varepsilon^{2/3}\log\varepsilon T}\right)\right)$

### v3

@@ 4.2 From the reduction to the accuracy lower bound
We use the following lower bound on inner product queries. Like a similar lower bound of Mir et al. [50], this theorem is proved using the reconstruction attacks of Dinur and Nissim [21] and Dwork et al. [23], together with the argument of De [17] that rules out reconstruction from the outputs of $(\varepsilon,\delta)$-DP algorithms.
Theorem 4.5 (Inner product queries lower bound (based on [21, 23, 50, 17])).
There are constants $c_{1}\geq 1$ and $c_{2}>0$ such that, for all sufficiently large $n>0$: if $\mathcal{A}$ is $\alpha$-accurate for $\mathsf{InnerProducts}_{k,n}$ (Definition 4.3) with $k=c_{1}n$ and $\alpha=c_{2}\sqrt{n}$, then $\mathcal{A}$ is not $(1,\frac{1}{3})$-differentially private.
We first prove Theorem 1.6 for $\varepsilon=1$ and then boost it to arbitrary $\varepsilon\leq 1$ using the reduction in Theorem A.1.
@@
For all $\delta\in(0,1)$, sufficiently large $w,T\in\mathbb{N}$ such that $w\leq T$, and all $(1,\delta)$-event-level-DP mechanisms that are $\alpha$-accurate for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$, if $\delta=o(\frac{1}{T})$, then
$\alpha=\Omega(\mathsf{min}(\sqrt{w},T^{1/4})).$
@@
Fix sufficiently large $w$ such that $w\leq\sqrt{T}$. Let $c_{1}\geq 1$ and $c_{2}>0$ be the constants from Theorem 4.5. Assume that $\mathcal{M}$ is a $\left(1,o(\frac{1}{T})\right)$-event-level-DP, $(\frac{c_{2}}{2}\sqrt{w})$-accurate mechanism for $\mathsf{CountDistinct}$ for turnstile streams of length $T$ with maximum flippancy at most $w$. Set $k=\frac{w}{2}$ and $n=\frac{k}{c_{1}}=\frac{w}{2c_{1}}$. This choice of $k$ and $n$ satisfies the conditions of Lemma 4.4 since the flippancy of the stream is at most $w=2k$. Moreover, for $w\leq\sqrt{T}$ we have that $(2k+1)n=(w+1)\frac{w}{2c_{1}}\leq w^{2}\leq T$. Therefore, $\mathcal{A}$ (Algorithm 4) is $\left(1,o(\frac{1}{T})\right)$-DP and $(c_{2}\sqrt{w})$-accurate for $\mathsf{InnerProducts}_{k,n}$. Since $\frac{1}{T}\leq\frac{1}{n}$ and $w=O(n)$, we get that $\mathcal{A}$ is $(1,o(\frac{1}{n}))$-DP and $c_{2}\sqrt{n}$-accurate for $\mathsf{InnerProducts}_{k,n}$, where $k=c_{1}n$.
@@
Finally, we invoke the reduction in Theorem A.1 to improve the dependence on $\varepsilon$ and complete the proof of Theorem 1.6.
@@
Suppose $\varepsilon<\frac{2}{T}$. For these values of $\varepsilon$, we prove an error lower bound of $\Omega(T)$ via a group privacy argument. Suppose for the sake of contradiction that $\alpha\leq T/4$. Consider universe $\mathcal{U}=[T]$. Let $x=\bot^{T}$ and $x^{\prime}$ be a stream of length $T$ such that $x[t]=t$ for all $t\in[T]$. These data streams differ in $T$ entries. Let $r[T]$ and $r^{\prime}[T]$ be the final outputs of $\mathcal{M}$ on input streams $x$ and $x^{\prime}$, respectively. By the accuracy of $\mathcal{M}$, we have $\Pr[r[T]\leq T/4]\geq 0.99.$ Applying Lemma 2.3 on group privacy with $\varepsilon\leq{2}/{T}$ and group size $\ell=T$, we get $\Pr[r^{\prime}[T]>T/4]\leq e^{2}\cdot\Pr[r[T]>T/4]+\frac{2\delta}{\varepsilon}\leq e^{2}\cdot 0.01+o(\frac{1}{T})<0.99$ for sufficiently large $T$. But $\mathsf{CountDistinct}(x^{\prime})=T$, so $\mathcal{M}$ is not $T/4$-accurate for $x^{\prime}$, a contradiction. Hence, $\alpha=\Omega(T)$.
Next, suppose $\varepsilon\geq\frac{2}{T}$. Lemma 4.6 provides a lower bound of $\alpha^{\prime}=\Omega\Big{(}\mathsf{min}\left(\sqrt{w},T^{1/4}\right)\Big{)}$ for ($\varepsilon^{\prime}=1$,$\delta^{\prime}=o(1/T)$)-event-level-DP, $\alpha^{\prime}$-accurate mechanisms for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$. By invoking Theorem A.1, we obtain the following lower bound on accuracy for $(\varepsilon,\delta)$-DP mechanisms where $\delta=\frac{\delta^{\prime}\varepsilon}{2}=o(\frac{\varepsilon}{T})$:
@@ 5 Item-level privacy lower bound
In this section, we prove Theorem 1.7 that provides strong lower bounds on the accuracy of any item-level differentially private mechanism for $\mathsf{CountDistinct}$ in the continual release model for turnstile streams. This lower bound is parameterized by $w$, the maximum flippancy of the input stream.
@@ 5.1 Reduction from $\mathrm{\sf Marginals}$
To prove our lower bounds for $\mathsf{CountDistinct}$, we reduce from the problem of approximating marginals in the batch model.
Definition 5.1 (Marginals).
The function $\mathrm{\sf Marginals}_{n,d}:\{0,1\}^{n\times d}\to[0,1]^{d}$ maps a dataset $y$ of $n$ records and $d$ attributes to a vector $(q_{1}(y),\dots,q_{d}(y)),$ where $q_{j}$, called the $j^{th}$ marginal, is defined as $q_{j}(y)=\frac{1}{n}\sum_{i=1}^{n}y[i][j].$
The reduction is presented in Algorithm 5. The privacy and accuracy guarantees of our reduction are stated in Lemma 5.3. In Section 5.2, we use Lemma 5.3 to complete the proof of Theorem 1.7.
Overview of the reduction.
Let $\mathcal{M}$ be an $(\varepsilon,\delta)$-DP and $\alpha$-accurate mechanism for $\mathsf{CountDistinct}$ in the continual release model. We use $\mathcal{M}$ to construct a $(O(\varepsilon),O(\delta))$-DP batch algorithm $\mathcal{A}$ that is $(\frac{\alpha}{n})$-accurate for $\mathrm{\sf Marginals}_{n,d}$. Consider a universe $\mathcal{U}=[n]\cup\{\perp\}$ for $\mathsf{CountDistinct}:\mathcal{U}_{\pm}^{T}\to\mathbb{N}$. The main idea in the construction (presented in Algorithm 5) is to force $\mathcal{M}$ to output an estimate of the marginals, one attribute at a time. Given a dataset $y\in\{0,1\}^{n\times d}$, the estimation of each marginal proceeds in two phases:
  • In phase one, $\mathcal{A}$ sends element $i$ to $\mathcal{M}$ for each record $y[i]$ with a $1$ in the first attribute. The answer produced by $\mathcal{M}$ at the end of phase one is an estimate of the sum of the first attribute of all records $y[1],\dots,y[n]$. This can be divided by $n$ to estimate the first marginal.
  • In phase two, $\mathcal{A}$ ‘clears the slate’ by sending $-i$ to $\mathcal{M}$ for each $y[i]$ with a $1$ in the first attribute.
It repeats this for each attribute, collecting the answers from $\mathcal{M},$ and then outputs its estimates for the marginals. In actuality, in both phases of estimating the $j^{\text{th}}$ marginal, $\mathcal{A}$ inputs $\perp$ for each $y[i]$ that has a 00 in the $j^{\text{th}}$ attribute. This algorithm is $(O(\varepsilon),O(\delta))$-DP for $\mathrm{\sf Marginals}_{n,d}$ since changing one record $y[i]$ in the input to the algorithm $\mathcal{A}$ will only change occurrences of a single element $i$ (to some other element $j$) in the input to the mechanism $\mathcal{M}$. Additionally, note that this reduction works equally well in the likes model where items can only be inserted when absent and deleted when present, since the stream produced in the reduction has this structure (see Theorem B.3).
@@
Let $\mathcal{A}$ be Algorithm 5. For all $\varepsilon\in(0,1],\delta\geq 0,\alpha\in\mathbb{R}^{+}$ and $d,n,w,T\in\mathbb{N}$, where $T\geq 2dn$ and $w\geq 2d$, if mechanism $\mathcal{M}$ is $(\varepsilon,\delta)$-item-level-DP and $\alpha$-accurate for $\mathsf{CountDistinct}$ for streams of length $T$ with maximum flippancy at most $w$ in the continual release model, then batch algorithm $\mathcal{A}$ is $(2\varepsilon,4\delta)$-DP and $\frac{\alpha}{n}$-accurate for $\mathrm{\sf Marginals}_{n,d}$.
@@
We start by reasoning about privacy. Fix neighboring datasets $y$ and $y^{\prime}$ that are inputs to batch algorithm $\mathcal{A}$ (Algorithm 5). (Datasets $y$ and $y^{\prime}$ differ in one row.) Let $x$ and $x^{\prime}$ be the streams constructed in Step 8 of $\mathcal{A}$ when it is run on $y$ and $y^{\prime}$, respectively. We show that $x$ and $x^{\prime}$ are $2$-item-neighbors. Suppose (w.l.o.g.) that $y$ and $y^{\prime}$ differ in the first row. Then $x$ and $x^{\prime}$ can only differ in the entries $z^{(j)}[1]$, where $j\in[d]$. The stream $x^{\prime}$ can be obtained from $x$ in a sequence of two item-level changes: first, replace all entries $z^{(j)}[1]$ for $j\in[d]$ with $\perp$, then replace a subset of these entries with the appropriate insertions and deletions, following phase one and two of Algorithm 5, to form $x^{\prime}$. Thus $x$ and $x^{\prime}$ are $2$-item-neighbors. Since $\mathcal{M}$ is $(\varepsilon,\delta)$-item-level-DP, and $\mathcal{A}$ only post-processes the outputs received from $\mathcal{M}$, closure under post-processing (Lemma 2.8) and group privacy (Lemma 2.3) imply that $\mathcal{A}$ is $(2\varepsilon,4\delta)$-DP (where we use the fact that $\frac{e^{2\varepsilon}-1}{e^{\varepsilon}-1}\leq 4$ for all $\varepsilon\in(0,1]$).
@@
Notice that: (1) The coins of $\mathcal{A}$ are the same as the coins of $\mathcal{M}$ (since the transformation from $\mathcal{M}$ to $\mathcal{A}$ is deterministic). (2) The marginals are computed in Step 10 of Algorithm 5 using the relationship described by Equation (7). (3) The maximum flippancy of the stream constructed in Algorithm 5 is at most $2d$, since each item $i\in\mathcal{U}$ is inserted and deleted at most once in each $z^{(j)}$ for $j\in[d]$. We obtain that $\mathcal{A}$ inherits its probability of success from $\mathcal{M}$:
@@ 5.2 From the reduction to the accuracy lower bound
For all $\varepsilon\in(0,2]$ , $\delta\in[0,1]$, $\gamma\in(0,1)$, $d,n\in\mathbb{N}$, and algorithms that are $(\varepsilon,\delta)$-differentially private and $\gamma$-accurate for $\mathrm{\sf Marginals}_{n,d}$, the following statements hold.
@@
To prove Theorem 1.7, we show that the lower bound holds for $\varepsilon=1$, and use Theorem A.1 to extend it to all $\varepsilon<1$. The approximate-DP lower bound (on the error term $\alpha$) in Theorem 1.7 is the minimum of two terms. To prove this bound, we need to establish that, for every possible range of parameters, at least one term serves as a lower bound for $\alpha$.
@@
For all $\delta\in(0,1]$, sufficiently large $w,T\in\mathbb{N}$ such that $w\leq T$, and all $(1,\delta)$-item-level-DP mechanisms that are $\alpha$-accurate for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$,
1 If $\delta>0$ and $\delta=o(1/T)$, then $\alpha=\Omega\Big{(}\mathsf{min}\left(\frac{\sqrt{w}}{\log w},\frac{T^{1/3}}{\log T}\right)\Big{)}$.
2 If $\delta=0$, then $\alpha=\Omega\Big{(}\mathsf{min}\left(w,\sqrt{T}\right)\Big{)}$.
@@
Let $\mathcal{A}$ be the algorithm for $\mathrm{\sf Marginals}_{n,d}$ with black-box access to an $\alpha$-accurate mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$, as defined in Algorithm 5. If $T\geq 2dn$ and $w\geq 2d$, then by Lemma 5.3, algorithm $\mathcal{A}$ is $(2,4\delta)$-differentially private and $(\frac{\alpha}{n})$-accurate for $\mathrm{\sf Marginals}_{n,d}$. We use Lemma 5.4 to lower bound $\alpha$.
Case 1 (Approximate DP, $\delta>0$, $\delta=o(1/n)$) :
@@
Now suppose $w=T^{2/3}$. The above argument gives a lower bound of $\Omega\left(\frac{\sqrt{T^{2/3}}}{\log T^{2/3}}\right)$ on the accuracy of $\mathcal{M}$. This lower bound applies to all $w>T^{2/3}$, since a mechanism that is $\alpha$-accurate for streams with maximum flippancy at most $w>w^{\prime}$ is also $\alpha$-accurate for streams with maximum flippancy at most $w^{\prime}$.
Case 2 (Pure DP, $\delta=0$):
The proof for $\delta=0$ is similar, except that we consider the cases $w\leq\sqrt{T}$ and $w>\sqrt{T}$ and use Item 2 from Lemma 5.4 instead of Item 1: Suppose $w\leq\sqrt{T}$. Pick a dimension $d=w/2$, and number of entries $n=\frac{T}{w}$. If $\frac{\alpha}{n}<1$, then by Lemma 5.3 and Item 1 of Lemma 5.4, $n=\Omega\left(\frac{n\cdot d}{\alpha\cdot\varepsilon}\right)$ which means that $\alpha=\Omega\left(\frac{d}{\varepsilon}\right)=\Omega(w)$. Otherwise, if $\alpha\geq n$, then $\alpha\geq\frac{T}{w}\geq\sqrt{T}\geq w$.
@@
This concludes the proof of Lemma 5.5. ∎
Finally, we extend the lower bounds for $\varepsilon=1$ from Lemma 5.5 to the general case of $\varepsilon<1$ using Theorem A.1.
@@
Suppose $\varepsilon<\frac{2}{T}$. For these values of $\varepsilon$, we prove an error lower bound of $\Omega(T)$, via a group privacy argument that is exactly the same as in the event-level lower bound (we direct the reader to the proof of Theorem 1.6 for more details).
@@
Now suppose $\varepsilon\geq\frac{2}{T}$. For $\delta>0$, Lemma 5.5 provides a lower bound of $\alpha^{\prime}=\Omega\Big{(}\mathsf{min}\left(\frac{\sqrt{w}}{\log w},\frac{T^{1/3}}{\log T}\right)\Big{)}$ on accuracy for ($\varepsilon^{\prime}=1$,$\delta^{\prime}=o(1/T)$)-item-level-DP, $\alpha^{\prime}$-accurate mechanisms for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$. By invoking Theorem A.1, we can extend this to the following lower bound for $(\varepsilon,\delta)$-DP mechanisms where $\delta=\frac{\delta^{\prime}\varepsilon}{2}=o(\frac{\varepsilon}{T})$:
@@
$\alpha=\frac{1}{\varepsilon}\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\log w},\frac{(\varepsilon T)^{1/3}}{\log\varepsilon T}\right)\right)=\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon\log w},\frac{T^{1/3}}{\varepsilon^{2/3}\log\varepsilon T}\right)\right).$